OpenStudy (anonymous):

series of (-1)* 3n^2+1/5n+2

OpenStudy (anonymous):

\[ \sum_{1}^{\infty}(-1)^n/\sqrt[3]{n} is it converges or diverges\]

OpenStudy (bahrom7893):

@Zarkon someone needs u :)

OpenStudy (rogue):

Is this your series?\[\sum_{n=0}^{\infty} (-1)^n \frac {3n^2 + 1}{5n + 2}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

is it diverges or converges

OpenStudy (rogue):

It diverges

OpenStudy (anonymous):

by which test

OpenStudy (rogue):

Dunno the name of the test, but since its an alternating series in the form of \[\sum_{n=0}^{\infty} (-1)^n a_n\]First thing you have to show for it to converge is that \[\lim_{n \rightarrow \infty} a_n = 0\]But for this, the limit is infinity, so the thing diverges just from that.

OpenStudy (anonymous):

i got it thank am from reno where r u from

OpenStudy (rogue):

np, I'm from nyc =)

OpenStudy (rogue):

nyc ftw =)

OpenStudy (anonymous):

\[\sum_{1}^{\infty}(-1)^n/\sqrt[3]{n}\]

OpenStudy (rogue):

converges

OpenStudy (anonymous):

which test

OpenStudy (rogue):

its a harmonic with p = 1.5. Since p >1, it converges, the alternation makes it converge even faster.

OpenStudy (zarkon):

it is just called the divergence test.

OpenStudy (anonymous):

oh i see

OpenStudy (zarkon):

the second one is the alterneting series test.

OpenStudy (anonymous):

is it conditionally

OpenStudy (rogue):

no, its absolute. There are no x terms for it to be conditional convergence.

OpenStudy (zarkon):

this \[\sum_{n=1}^{\infty}\frac{1}{\sqrt[3]{n}}\] diverges but \[\sum_{n=1}^{\infty}\frac{(-1)^n}{\sqrt[3]{n}}\] converges by the AST

OpenStudy (rogue):

:o the first diverges? didn't know.

OpenStudy (zarkon):

yes...p-series test.

OpenStudy (rogue):

but p = 1.5, so...

OpenStudy (zarkon):

p=1/3

OpenStudy (anonymous):

so its only p-series and ABS conv noALT

OpenStudy (rogue):

I thought it was \[\sum_{n = 0}^{\infty} \frac {1}{n^P}\]

OpenStudy (zarkon):

\[\sum_{n=1}^{\infty}\frac{1}{\sqrt[3]{n}}=\sum_{n=1}^{\infty}\frac{1}{n^{1/3}}\]

OpenStudy (rogue):

oh, right, what am I thinking :3

OpenStudy (zarkon):

you cant start at zero either

OpenStudy (anonymous):

SO I THINK IS CONV p-series and ALT series right

OpenStudy (rogue):

thanks, my mind is so wacky today ;/ alternating series test.

OpenStudy (zarkon):

the series is conditionally convergent

OpenStudy (anonymous):

why is conditionally cqan u tel me so i can understanfd

OpenStudy (zarkon):

\[\sum_{n=1}^{\infty}\frac{(-1)^n}{\sqrt[3]{n}}\] converges by the AST \[\sum_{n=1}^{\infty}\left|\frac{(-1)^n}{\sqrt[3]{n}}\right|=\sum_{n=1}^{\infty}\frac{1}{\sqrt[3]{n}}\] diverges by the p-series test. therefore it is conditionally convergent

OpenStudy (anonymous):

\[x \cos (x) what is the series\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

is that \[\sum_{0}^{\infty}(-1)^n/(2n_1)!*x^2n+1\]

OpenStudy (anonymous):

is that right

OpenStudy (zarkon):

I think I know what you wrote..latex looks bad...I believe it is ok

OpenStudy (anonymous):

oh that was 2n+1

OpenStudy (anonymous):

can u corct it plz zarkon

OpenStudy (zarkon):

\[\sum_{n=0}^{\infty}\frac{(-1)^n}{(2n)!}x^{2n+1}\]

OpenStudy (anonymous):

this look like e^x series

OpenStudy (anonymous):

but anm try to wirte xcosx series

OpenStudy (anonymous):

is there anyone can help

OpenStudy (anonymous):

r u there zarkon

OpenStudy (rogue):

The Maclauren series for cosine is\[\cos x = 1 - \frac {x^2}{2!} + \frac {x^4}{4!} - \frac {x^6}{6!} + ...\]So when you multiply cosine x by x, your series is \[x \cos x = x \sum_{n=0}^{\infty} (-1)^n \frac {x^{2n}}{(2n)!} = \sum_{n=0}^{\infty} (-1)^n \frac {x^{2n+1}}{(2n)!}\]

OpenStudy (anonymous):

thanks